# Physics-Driven Construction of Compact Primitive Gaussian Density Fitting Basis Sets

### v1

@@ 4.1 Training MADF model parameters
The MADF model parameters were tuned to ensure that that the DF errors for the respective training sets did not exceed target accuracies of ±$\pm 20\mu E_{h}$ and ±$\pm 10\mu E_{h}$ per electron for HF and MP2 energies, respectively. This target range for errors was determined by analyzing the DF errors of manually-optimized DFBSs; the majority have larger errors than these targets (see Section 4.2.1).
4.1.1 Training $\tau_{\text{H}}$
The subset of all hydrogen-containing molecules from the TS1 set (Section 3) was used to determine the recommended value of $\tau_{\text{H}}$. While varying $\tau_{\text{H}}$ the remaining MADF parameters were fixed ($\tau_{1}=\tau_{2}=10^{-8},\zeta=1.2$) sufficiently closely to their asymptotic limits ($\tau\to 0$, $\zeta\to 1$) without causing ill-conditioning and excessive costs. Figures 2(a) and 2(b) show the DF errors of HF and MP2 energies, respectively, with $\tau_{\text{H}}$ varying between $10^{-8}$ and $10^{-4}$. The errors are essentially constant with $\tau_{\text{H}}\leq 10^{-6}$, thus $\tau_{\text{H}}=10^{-6}$ is recommended. Since hydrogen has almost negligible relativistic effects, this value will be used henceforth for both relativistic and nonrelativistic calculations.
Figure 2: Variation of DF errors of nonrelativistic (a) HF and (b) MP2 energies of the H-containing subset of the TS1 training set vs the $\tau_{\text{H}}$ model parameter of the MADF generator.
@@ 4.1.2 Training $\tau_{1,2}$ and $\zeta$ for nonrelativistic OBS
The variation of DF errors with $\tau_{2}$ was studied next, by keeping the rest of the parameters fixed ($\tau_{1}=10^{-8},\zeta=1.2$) sufficiently closely to their asymptotic limits ($\tau\to 0$, $\zeta\to 1$) without causing ill-conditioning and excessive costs. As Figures 3(a) and 3(b) indicate, the DF errors of HF and MP2 are largely converged at $\tau_{2}=10^{-5}$. Next, $\tau_{2}$ was fixed at $10^{-5}$ and $\tau_{1}$ was varied (Figures 4(a) and 4(b)). The DF errors are sufficiently converged with $\tau_{1}=10^{-6}$. Finally, with $\tau_{1}=10^{-6}$ and $\tau_{2}=10^{-5}$ variation of DF errors with $\zeta$ was studied (Figures 5(a) and 5(b)). For $\zeta>1.4$ the max DF errors increase due to the insufficient spanning of the OBS AO product space by the DFBS AOs, whereas for $\zeta<1.2$ (not shown) the onset of ill-conditioning in DFBS makes the MADF model not numerically stable. $\zeta=1.4$ ensures sufficient convergence of the DF errors without undue numerical problems.
Note that $\zeta=1.4$ is significantly smaller than the exponent ratio 1.8 used by the AutoAux generator ( 1.8)61 and ratio 2 used by PySCF’s generator65 for even-tempered DFBS construction. Our findings confirm that the relatively small exponent ratios are indeed necessary for accurate density fitting. For example, exponent ratios in the $[1.3,1.5]$ range are quite common in the $L\leq 2$ channels of the manually-optimized def2-QZVPP-RIFIT DFBS (see Figure 1). Clearly even-tempered spanning of the AO product space is suboptimal for practical basis sets.
@@
Parameters $\tau_{1,2}$ and $\zeta$ were trained again using the relativistic HF and MP2 energies of the TS2 set. $\tau_{2}$ was varied first, with $\tau_{1}=10^{-8}$ and $\zeta=1.2$ (Figures 6(a) and 6(b)); just as in nonrelativistic computations $\tau_{2}=10^{-5}$ ensures sufficient convergence of the DF errors. $\tau_{1}$ was then varied with fixed $\tau_{2}=10^{-5}$ and $\zeta=1.2$ (Figures 7(a) and 7(b)). In contrast to the nonrelativistic case, a much smaller $\tau_{1}$ is needed for sufficient convergence of the DF errors. The variation of relativistic DF errors with $\zeta$ (Figures 8(a) and 8(b)) is similar to the nonrelativistic case, with $\zeta=1.4$ deemed sufficient. Yet again, this is a much smaller ratio than the ratios used by even-tempered DFBS generators.61, 65
@@ 5 Summary
This work introduced the MADF algorithm for generating density-fitting basis set composed of primitive solid-harmonic Gaussian AOs and suited for nonrelativistic and relativistic (all-electron) computations with mean-field and correlated electronic structure models. Not using even-tempered AO sets for spanning the AO product space, as in other comparable generators,61, 65 produces compact and accurate candidate pools of DFBS AOs while sufficiently reducing the numerical redundancy. The subsequent pruning of the candidate DFBS pool based on the 2-body energy as the importance metric allows to naturally reduce the relevant range of angular momentum channels and exponent spans with minimal use of heuristics. Generation of primitive (rather than contracted) DF AOs allows to keep the integral evaluation costs optimally low. Using only 4 model parameters MADF generates basis sets that match or exceed the accuracy-to-size ratio of the state-of-the-art DFBS generators; the comparison is particularly favorable with OBSs suited for relativistic all-electron computations. Computations utilizing MADF DFBS, of course, can also benefit from system-specific DFBS compressions.67, 99
Although picture change and other relativistic effects on the 2-particle interaction were not considered here, it will be interesting to examine the impact of such effects on the optimal MADF model parameters in the future; it is known that there are nontrivial differences in the requirements on the DFBS that stem from relativistic contributions to the effective 2-particle interactions.13 Even in the nonrelativistic framework much more testing of the approach will be clearly needed. Most important is to test the performance in the context of higher-order correlated methods (especially coupled cluster), for DF approximation of non-Coulomb operators, for properties other than energy, and for other basis set families (e.g., multiply-augmented correlation-consistent sets, ANO, etc.). Early experiences with MADF DFBSs appear promising. For example, we recently utilized the MADF DFBS generator for explicitly-correlated (F12) coupled-cluster computations with aug-cc-pV7Z OBS (for which no manually-optimized DFBS is available) where we found that the DF errors with the MADF DFBS are significantly smaller than the use of the manually-optimized aug-cc-pV6Z-RIFIT basis. These and other experiences with MADF DFBSs will be reported elsewhere.
@@
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, and Office of Basic Energy Sciences, Scientific Discovery through the Advanced Computing (SciDAC) program under Award DE-SC0022263. This project used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract DE-AC02-05CH11231 using NERSC award ERCAP-0024336. The authors acknowledge Advanced Research Computing at Virginia Tech (https://arc.vt.edu/) for providing computational resources and technical support that have contributed to the results reported within this paper. The development of the Libint software library is supported by the Office of Advanced Cyberinfrastructure, National Science Foundation (Award OAC-2103738).
@@ Appendix A Superposition of atomic densities (SOAD)
Unlike the SAD method,85 with computes the Fock matrix from the atomic (optionally, spin-free) ensemble density matrix computed in OBS self-consistently, SOAD Fock matrix is computed from fixed spin-free ensemble occupancies ${\bf n}_{\mathrm{SOAD}}$ of the minimal basis set (MBS) AOs, without self-consistency. ${\bf n}_{\mathrm{SOAD}}$ for a single atom in an orthonormal MBS is obtained by smearing each subshell’s electrons evenly among all of its orbitals; e.g., for a carbon atom occupancies of 1s, 2s, 2px, 2py, and 2pz MBS AOs are 2, 2, 2/3, 2/3, and 2/3, respectively. This recipe has been referred to as average-of-configuration (AOC)100 and canonical ensemble101. The number of electrons in each subshell is defined by the electron configuration of that atom’s neutral ground state taken from NIST Atomic Spectra Database102. For a molecule ${\bf n}_{\mathrm{SOAD}}$ is obtained by concatenation of the atomic occupancies, without taking into account the effects of nonunit molecular MBS metric. In this work the MINI basis103, 104 was used as MBS for $Z\leq 54$, and the ANO-RCC-MB basis105 for $54<Z\leq 96$.
${\bf n}_{\mathrm{SOAD}}$ is used to construct SOAD Fock matrix in OBS AO basis, $\mathbf{F_{\mathrm{SOAD}}}$, as follows:
@@
where $h^{\mu}_{\nu}$ are the matrix elements of the nonrelativistic 1-body (core) Hamiltonian, $m$ runs over MBS AOs and $\mu,\nu$ over OBS AOs.

### v2

@@ 4.1 Training MADF model parameters
The MADF model parameters were tuned to ensure that that the DF errors for the respective training sets did not exceed target accuracies of $\pm 20\mu E_{h}$ and $\pm 10\mu E_{h}$ per electron for HF and MP2 energies, respectively. This target range for errors was determined by analyzing the DF errors of manually-optimized DFBSs; the majority have larger errors than these targets (see Section 4.2.1).
@@ 4.1.1 Training $\tau_{1,2}$ and $\zeta$ for nonrelativistic OBS
The variation of DF errors with $\tau_{2}$ was studied next, by keeping the rest of the parameters fixed ($\tau_{1}=10^{-8},\zeta=1.2$) sufficiently closely to their asymptotic limits ($\tau\to 0$, $\zeta\to 1$) without causing ill-conditioning and excessive costs. As Figures 2(a) and 2(b) indicate, the DF errors of HF and MP2 are largely converged at $\tau_{2}=10^{-5}$. Next, $\tau_{2}$ was fixed at $10^{-5}$ and $\tau_{1}$ was varied (Figures 3(a) and 3(b)). The DF errors are sufficiently converged with $\tau_{1}=10^{-6}$. Finally, with $\tau_{1}=10^{-6}$ and $\tau_{2}=10^{-5}$ variation of DF errors with $\zeta$ was studied (Figures 4(a) and 4(b)). For $\zeta>1.4$ the max DF errors increase due to the insufficient spanning of the OBS AO product space by the DFBS AOs, whereas for $\zeta<1.2$ (not shown) the onset of ill-conditioning in DFBS makes the MADF model not numerically stable. $\zeta=1.4$ ensures sufficient convergence of the DF errors without undue numerical problems.
Note that $\zeta=1.4$ is significantly smaller than the exponent ratio 1.8 used by the AutoAux generator ( 1.8)61 and ratio 2 used by PySCF’s generator67 for even-tempered DFBS construction. Our findings confirm that the relatively small exponent ratios are indeed necessary for accurate density fitting. For example, exponent ratios in the $[1.3,1.5]$ range are quite common in the $L\leq 2$ channels of the manually-optimized def2-QZVPP-RIFIT DFBS (see Figure 1). Clearly even-tempered spanning of the AO product space is suboptimal for practical basis sets.
@@
Parameters $\tau_{1,2}$ and $\zeta$ were trained again using the relativistic HF and MP2 energies of the TS2 set. $\tau_{2}$ was varied first, with $\tau_{1}=10^{-8}$ and $\zeta=1.3$. Here we start with $\zeta=1.3$ instead of $\zeta=1.2$ like in the non-relativistic case because $\zeta=1.2$ is too small and causes numerical issues with the relativistic OBSs. (Figures 5(a) and 5(b)); just as in nonrelativistic computations $\tau_{2}=10^{-5}$ ensures sufficient convergence of the DF errors. $\tau_{1}$ was then varied with fixed $\tau_{2}=10^{-5}$ and $\zeta=1.3$ (Figures 6(a) and 6(b)). In contrast to the nonrelativistic case, a much smaller $\tau_{1}$ is needed for sufficient convergence of the DF errors. The variation of relativistic DF errors with $\zeta$ (Figures 7(a) and 7(b)) is similar to the nonrelativistic case, with $\zeta=1.4$ deemed sufficient. Yet again, this is a much smaller ratio than the ratios used by even-tempered DFBS generators.61, 67
@@ 4.2.1 The G2 set
We have also compared the performance of primitive DFBSs produced by MADF with contracted DFBSs produced by acCD. From Figure 9(a) it is clear that MADF DFBSs produce smaller DF errors in HF energies compared to acCD DFBSs. As can be seen in Figure 9(b) the MP2 errors for cc-pVDZ are smaller with MADF compared to acCD and for cc-pVTZ and cc-pVQZ the magnitude MP2 errors are $5\mu E_{h}$ per electron, which is sufficient. MADF also provides significantly smaller DFBSs than that of acCD with virtually similar errors for cc-pVTZ and cc-pVQZ as seen in Figure 9(c).
Figure 9: Comparison of DF errors of (a) HF and (b) MP2 energies of the closed-shell subset of the G2 test set obtained with acCD and MADF DFBSs. (c) Comparison of DFBS sizes, relative to that of the corresponding OBS. For acCD, the number of contracted DF AOs was used.
@@ 5 Summary
This work introduced the MADF algorithm for generating density-fitting basis set composed of primitive solid-harmonic Gaussian AOs and suited for nonrelativistic and relativistic (all-electron) computations with mean-field and correlated electronic structure models. Not using even-tempered AO sets for spanning the AO product space, as in other comparable generators,61, 67 produces compact and accurate candidate pools of DFBS AOs while sufficiently reducing the numerical redundancy. The subsequent pruning of the candidate DFBS pool based on the 2-body energy as the importance metric allows to naturally reduce the relevant range of angular momentum channels and exponent spans with minimal use of heuristics. Generation of primitive (rather than contracted) DF AOs allows to keep the integral evaluation costs optimally low. Using only 3 model parameters MADF generates basis sets that match or exceed the accuracy-to-size ratio of the state-of-the-art DFBS generators; the comparison is particularly favorable with OBSs suited for relativistic all-electron computations. Computations utilizing MADF DFBS, of course, can also benefit from system-specific DFBS compressions.70, 105
Although picture change and other relativistic effects on the 2-particle interaction were not considered here, it will be interesting to examine the impact of such effects on the optimal MADF model parameters in the future; it is known that there are nontrivial differences in the requirements on the DFBS that stem from relativistic contributions to the effective 2-particle interactions.13 Even in the nonrelativistic framework much more testing of the approach will be clearly needed. Most important is to test the performance in the context of higher-order correlated methods (especially coupled cluster), for DF approximation of non-Coulomb operators, for properties other than energy, and for other basis set families (e.g., multiply-augmented correlation-consistent sets, ANO, etc.). Early experiences with MADF DFBSs appear promising. For example, we recently utilized the MADF DFBS generator for explicitly-correlated (F12) coupled-cluster computations with aug-cc-pV7Z OBS (for which no manually-optimized DFBS is available) where we found that the DF errors with the MADF DFBS are significantly smaller than the use of the manually-optimized aug-cc-pV6Z-RIFIT basis.106 These and other experiences with MADF DFBSs will be reported elsewhere.
@@
Work by KAS was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, and Office of Basic Energy Sciences, Scientific Discovery through the Advanced Computing (SciDAC) program under Award DE-SC0022263. Work by EFV was supported by the US Department of Energy, Office of Science, via award DE-SC0022327. This project used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract DE-AC02-05CH11231 using NERSC award ERCAP-0024336. The authors acknowledge Advanced Research Computing at Virginia Tech (https://arc.vt.edu/) for providing computational resources and technical support that have contributed to the results reported within this paper. The development of the Libint software library is supported by the Office of Advanced Cyberinfrastructure, National Science Foundation (Award OAC-2103738).
@@ Appendix A Superposition of atomic densities (SOAD)
Unlike the SAD method,89 with computes the Fock matrix from the atomic (optionally, spin-free) ensemble density matrix computed in OBS self-consistently, SOAD Fock matrix is computed from fixed spin-free ensemble occupancies ${\bf n}_{\mathrm{SOAD}}$ of the minimal basis set (MBS) AOs, without self-consistency. ${\bf n}_{\mathrm{SOAD}}$ for a single atom in an orthonormal MBS is obtained by smearing each subshell’s electrons evenly among all of its orbitals; e.g., for a carbon atom occupancies of 1s, 2s, 2px, 2py, and 2pz MBS AOs are 2, 2, 2/3, 2/3, and 2/3, respectively. This recipe has been referred to as average-of-configuration (AOC)107 and canonical ensemble108. The number of electrons in each subshell is defined by the electron configuration of that atom’s neutral ground state taken from NIST Atomic Spectra Database109. For a molecule ${\bf n}_{\mathrm{SOAD}}$ is obtained by concatenation of the atomic occupancies, without taking into account the effects of nonunit molecular MBS metric. In this work the MINI basis110, 111 was used as MBS for $Z\leq 54$, and the ANO-RCC-MB basis112 for $54<Z\leq 96$. ${\bf n}_{\mathrm{SOAD}}$ is used to construct SOAD Fock matrix in OBS AO basis, $\mathbf{F_{\mathrm{SOAD}}}$, as follows:
@@
where $h^{\mu}_{\nu}$ are the matrix elements of the 1-body (core) Hamiltonian (sf-1eX2C90, 91, 92 in this work) , $m$ runs over MBS AOs and $\mu,\nu$ over OBS AOs. It should be noted that Eq. 26 is similar to the projection-free guess proposed by Norman and Jensen113 for Dirac-Hartree-Fock/Kohn-Sham.